Pompeii Pizza Club owns three identical restaurants popular for their specialty pizzas. Each restaurant has a debt-equity ratio of 30 percent and makes interest payments of $48,000 at the end of each year. The cost of the firm’s levered equity is 17 percent. Each store estimates that annual sales will be $1.365 million; annual cost of goods sold will be $775,000; and annual general and administrative costs will be $435,000. These cash flows are expected to remain the same forever. The corporate tax rate is 21 percent.

  

a.

Use the flow to equity approach to determine the value of the company’s equity. (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to 2 decimal places, e.g., 1,234,567.89)

b. What is the total value of the company? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to 2 decimal places, e.g., 1,234,567.89)
Answer #1

C D E Income Statement Sales 4095000 Cost of goods sold 2325000 G&A costs 1305000 EBIT 465000 Interest 144000 EBT 321000 Tax

EXCEL FORMULA:

Income Statement Sales Cost of goods sold G&A costs EBIT Interest EBT Tax Net income =1365000*3 =775000*3 =435000*3 C3-C4-C5